# Adaptive Cyclic Diffusion for Inference Scaling

### v1

@@ 5 Experiments
We evaluate our proposed method on a diverse suite of challenging tasks that require generating samples in sparse or previously unexplored regions at training phase: (1) toy Mixture of Gaussian(MoG), (2) point-mass maze navigation (park2024ogbench, ), (3) Sudoku puzzle completion (palm2018recurrentrelationalnetworks, ; wang2019satnetbridgingdeeplearning, ), (4) path generation on unseen pixel maze image (maze, ), and (5) molecular structure prediction (hoogeboom2022, ). These tasks present distinct challenges: MoG shows the importance of multiple go-back temperature pool as the proof-of-concept; point-mass maze requires long-horizon planning over 1000 steps; Sudoku demands logical consistency across row, column, and block constraints; path generation on unseen maze image tests generalization to novel environmental structures; and molecular structure prediction requires generating valid 3D conformations under chemical and physical constraints.
@@ 5.1 Baselines
We compare ABCD against several strong representative baselines: Base Diffusion: The standard diffusion model without additional computational scaling, serving as our primary baseline. Best-of-$N$ (BoN): A computationally scaled variant that generates $N$ independent samples and selects the best according to a task-specific verifier used consistently across all models. Diffusion Beam Search (BS) (oshima2025inferencetimetexttovideoalignmentdiffusion, ): Combines diffusion with beam search to scale the denoising process, pruning unpromising candidates during generation. Sequential Monte Carlo Diffusion (SMC) (singhal2025generalframeworkinferencetimescaling, ; wu2023practical, ): Employs resampling during the denoising process to focus computational resources on more promising trajectories. Implementation details were adapted from  singhal2025generalframeworkinferencetimescaling to suit our experimental setting. Search over Paths (SoP) (ma2025inferencetimescalingdiffusionmodels, ): Enhances generation by allowing limited backward steps (adding noise) to promising samples, enabling exploration of wider solution spaces compared to strictly unidirectional approaches.
All baselines used the same pre-trained diffusion model per tasks, and inference-time budgets were expanded as much as possible while preserving their core operational characteristics. In BS and SoP, inference-time compute was increased by more frequent expansion and selection, while SMC scaled compute via higher resampling frequency, and BoN by increasing the number of particles $N$. For fair comparison, the total number of particles was set equal across all methods (except BoN).
@@ 5.4 Path generation on unseen pixel maze image
Setup. We evaluated ABCD’s adaptability on a path-generation task using mazes of previously unseen sizes (maze, ). Every method leveraged the same diffusion model pretrained on small mazes (sizes 4, 5, and 6) and was then tested on larger, structurally distinct mazes zhang2025tscendtesttimescalablemctsenhanced . This setting rigorously probes each algorithm’s capacity to generalize to novel spatial configurations at inference time.
@@ 5.5 Molecule generation
Result. As shown in the Figure 4(Right), ABCD significantly outperforms all baselines by achieving higher molecular stability with less inference wall-clock time. For instance, ABCD reaches a molecule stability of 0.97 within approximately 20 seconds, a threshold that other baselines either fail to attain or require considerably more computational time to approach. Notably, our method achieved a peak molecule stability of approximately 0.99, which is about 5.32% higher than the peak stability of 0.94 achieved by SoP (ma2025inferencetimescalingdiffusionmodels, ), the next best-performing baseline. Furthermore, the performance gap between ABCD and the baselines widens at higher stability levels, emphasizing the robustness of our method in complex molecular generation tasks.
@@ 5.6 OGBench Point Maze
Result. As shown in Figure 6, ABCD consistently surpasses other baselines. Moreover, only ABCD achieves perfect performance on both mazes. It reflects the advantages of our method: by quickly denoising to $t=0$, it locates valid trajectories almost immediately, rather than spending compute on uniformly long reverse sampling path. In maze giant, the gap widens because larger mazes demand more on both global exploration (to reach the goal) and fine refinement (to thread tight corridors).
@@ 5.7 Do we need to care about go-back temperature level?
Instance-Level optimal temperature. Figure 7(a–b) plots cycle-wise accuracy for two Sudoku instances (14 vs. 44 givens) under fixed go-back noise levels. Certain noise settings plateau after only a few cycles and the plateau point differs between puzzles, indicating that each instance possesses its own optimal noise level.
@@ 5.8 Do we need adaptive terminal condition?
ABCD automatically assigns proper thinking time. Figure 7(d) demonstrates that ABCD allocates more thinking cycles to harder puzzles (fewer givens) and terminates earlier on simpler ones, validating that our adaptive terminal criterion scales compute with problem difficulty. Furthermore, the stop-flag hyperparameter $\kappa$ directly controls exploration depth: larger $\kappa$ values enforce additional cycles before stopping, which boosts success on challenging instances. Identical trends hold on the Pixel Maze task.

### v2

@@ 5 Experiments
We evaluate our proposed method on a diverse suite of challenging tasks that require generating samples in sparse or previously unexplored regions at training phase: (1) toy Mixture of Gaussian(MoG), (2) point-mass maze navigation (park2024ogbench, ), (3) Sudoku puzzle completion (palm2018recurrentrelationalnetworks, ; wang2019satnetbridgingdeeplearning, ), (4) path generation on unseen pixel maze image (maze, ), and (5) molecular structure prediction (hoogeboom2022, ). These tasks present distinct challenges: MoG shows the importance of multiple go-back temperature pool as the proof-of-concept; point-mass maze requires long-horizon planning over 1000 steps; Sudoku demands logical consistency across row, column, and block constraints; path generation on unseen maze image tests generalization to novel environmental structures; and molecular structure prediction requires generating valid 3D conformations under chemical and physical constraints. Detailed task configurations are discussed in Appendix B.2.
@@ 5.1 Baselines
We compare ABCD against several strong representative baselines: Base Diffusion: The standard diffusion model without additional computational scaling, serving as our primary baseline. Best-of-$N$ (BoN): A computationally scaled variant that generates $N$ independent samples and selects the best according to a task-specific verifier used consistently across all models. Diffusion Beam Search (BS) (oshima2025inferencetimetexttovideoalignmentdiffusion, ): Combines diffusion with beam search to scale the denoising process, pruning unpromising candidates during generation. Sequential Monte Carlo Diffusion (SMC) (singhal2025generalframeworkinferencetimescaling, ; wu2023practical, ): Employs resampling during the denoising process to focus computational resources on more promising trajectories. Implementation details were adapted from  singhal2025generalframeworkinferencetimescaling to suit our experimental setting. Search over Paths (SoP) (ma2025inferencetimescalingdiffusionmodels, ): Enhances generation by allowing limited backward steps (adding noise) to promising samples, enabling exploration of wider solution spaces compared to strictly unidirectional approaches. Detailed Baseline configurations are discussed in Appendix B.3.
All baselines used the same pre-trained diffusion model per tasks, and inference-time budgets were expanded as much as possible while preserving their core operational characteristics. In BS and SoP, inference-time compute was increased by more frequent expansion and selection, while SMC scaled compute via higher resampling frequency, and BoN by increasing the number of particles $N$. For fair comparison, the total number of particles was set equal across all methods (except BoN). Additional experimental details for each tasks are provided in the Appendix B.6B.8B.10B.12.
@@ 5.2 Mixture of Gaussian
Setup. We construct two toy datasets to illustrate the need for multiple go-back noise levels and terminal condition during inference refinement (Figure 2(a)). Dataset 1 has modes clustered locally, so small-step refinement suffices; Dataset 2’s modes lie in distant regions, requiring occasional large “jumps” to escape poor initial prediction. We train separate diffusion models on each and condition sampling on distance to a target (red star). Full details appear in Appendix B.4.
@@ 5.3 Sudoku puzzle completion
Setup. We evaluated ABCD against baseline methods on the Sudoku puzzle completion task (palm2018recurrentrelationalnetworks, ; wang2019satnetbridgingdeeplearning, ), a canonical logical-reasoning benchmark. In Sudoku, a 9×9 grid must be filled under row, column, and 3×3-block constraints, with each new grid filled based on the current grid state. An adaptive search strategy—broadly exploring when few clues are given and then locally refining once most entries are correct—proves particularly effective. We trained all models on puzzles with 31–42 given digits wang2019satnetbridgingdeeplearning and tested them on more challenging instances containing only 17–28 givens palm2018recurrentrelationalnetworks , thereby measuring each method’s ability to produce logically consistent solutions. See Appendix B.8 for additional experimental details.
@@ 5.4 Path generation on unseen pixel maze image
Setup. We evaluated ABCD’s adaptability on a path-generation task using mazes of previously unseen sizes (maze, ). Every method leveraged the same diffusion model pretrained on small mazes (sizes 4, 5, and 6) and was then tested on larger, structurally distinct mazes zhang2025tscendtesttimescalablemctsenhanced . This setting rigorously probes each algorithm’s capacity to generalize to novel spatial configurations at inference time. Additional experimental details are provided in Appendix B.10.
@@ 5.5 Molecule generation
Result. As shown in the Figure 4(Right), ABCD significantly outperforms all baselines by achieving higher molecular stability with less inference wall-clock time. For instance, ABCD reaches a molecule stability of 0.97 within approximately 20 seconds, a threshold that other baselines either fail to attain or require considerably more computational time to approach. Notably, our method achieved a peak molecule stability of approximately 0.99, which is about 5.32% higher than the peak stability of 0.94 achieved by SoP (ma2025inferencetimescalingdiffusionmodels, ), the next best-performing baseline. Furthermore, the performance gap between ABCD and the baselines widens at higher stability levels, emphasizing the robustness of our method in complex molecular generation tasks. See Appendix B.12 for more details.
@@ 5.6 OGBench Point Maze
Result. As shown in Figure 6, ABCD consistently surpasses other baselines. Moreover, only ABCD achieves perfect performance on both mazes. It reflects the advantages of our method: by quickly denoising to $t=0$, it locates valid trajectories almost immediately, rather than spending compute on uniformly long reverse sampling path. In maze giant, the gap widens because larger mazes demand more on both global exploration (to reach the goal) and fine refinement (to thread tight corridors). Further results analysis and visualization of the trajectories can be founded in the Appendix B.7.
@@ 5.7 Do we need to care about go-back temperature level?
Instance-Level optimal temperature. Figure 7(a–b) plots cycle-wise accuracy for two Sudoku instances (14 vs. 44 givens) under fixed go-back noise levels. Certain noise settings plateau after only a few cycles and the plateau point differs between puzzles, indicating that each instance possesses its own optimal noise level. (see Appendix C.1).
@@ 5.8 Do we need adaptive terminal condition?
ABCD automatically assigns proper thinking time. Figure 7(d) demonstrates that ABCD allocates more thinking cycles to harder puzzles (fewer givens) and terminates earlier on simpler ones, validating that our adaptive terminal criterion scales compute with problem difficulty. Furthermore, the stop-flag hyperparameter $\kappa$ directly controls exploration depth: larger $\kappa$ values enforce additional cycles before stopping, which boosts success on challenging instances. Identical trends hold on the Pixel Maze task. (see Appendix B.11).